OpenStudy (anonymous):

A cone is inscribed in a sphere with a diameter of 6. What is the maximum possible volume of the cone... very hard problem

OpenStudy (anonymous):

im in calc BC by the way did you get 32pi/3?

OpenStudy (zehanz):

See image. basecircle of cone has radius 6sin(2 theta) height of cone = 6 + 6cos(2 theta) V(theta)=1/3 * pi * (6sin(2*theta))²(6+6cos(2 theta)) Now calculate the derivative and solve V'= 0...

OpenStudy (anonymous):

Wait... its not necessarily an equilateral

OpenStudy (zehanz):

In my drawing theta = alpha. The base of the cone (DE) is not equal to DC or DE, but DC=DE, so u can use my formula to calculate the optimum volume. Going to sleep now...Happy New Year! (already in Europe)

OpenStudy (zehanz):

\[V(\theta)=\frac{ 1 }{ 3 }\pi r^2h=\frac{ 1 }{ 3 }\pi \cdot 36\sin^22\theta(6+6\cos2\theta)\]\[V(\theta)=72\pi \sin^22\theta(1+\cos2\theta)\]

OpenStudy (sirm3d):

maximum volume is \(\displaystyle \frac{32\pi}{3}\)

OpenStudy (zehanz):

Derivative of V:\[V'(\theta)=72\pi2\sin2\theta \cdot 2\cos2\theta(1+\cos2\theta)+72\pi \sin^22\theta \cdot-2\sin2\theta=0\] Divide everything by 144pi sin2θ: (because 2θ > 0, sin2θ >0, so dividing is OK)\[2\cos2\theta(1+\cos2\theta)-\sin^22\theta=0\]Use \[\cos^2x+\sin^2x=1\]to convert sin²2θ:\[2\cos2\theta+2\cos^22\theta+\cos^22\theta-1=0\]\[3\cos^22\theta+2\cos2\theta-1=0\]abc-formula:\[\cos2\theta=\frac{ -2 \pm \sqrt{4+12} }{ 6 }=-\frac{ 1 }{ 3 }\pm \frac{ 2 }{ 3 }\]so\[\cos2\theta=-1 \vee \cos2\theta=\frac{ 1 }{ 3 }\]Now, the first answer gives 2θ=π, which would mean a top angle op 180 degrees. This "cone" would have a volume of zero, so we have a minimum (not interesting).

OpenStudy (zehanz):

The second answer, \[\cos 2\theta = \frac{ 1 }{ 3 }\]would mean\[\cos^22\theta=\frac{ 1 }{ 9 }\]and thus\[\sin^22\theta=\frac{ 8 }{ 9 }\]so\[\sin2\theta=\pm \frac{ 8 }{ 9 }=\pm \frac{ 2 }{ 3 }\sqrt{2}\]Of course, because\[0<2\theta<\pi\]after all, it has to be a cone, we see that \[\sin2\theta=\frac{ 2 }{ 3 }\sqrt{2}\]We can now use these values in the formula for the volume. Maximum volume is:\[V=72\pi \frac{ 8 }{ 9 }(1+\frac{ 1 }{ 3 })=64\pi \frac{ 4 }{ 3 }=\frac{ 256\pi }{ 3 }\] @sirm3d: my answer is different from yours!

OpenStudy (zehanz):

Typo: \[\sin2\theta=\pm \sqrt{\frac{ 8 }{ 9 }}=\pm \frac{ 2 }{ 3 }\sqrt{2}\]